A glass surface is coated by an oil film of uniform thickness 1.00 × 10-4 cm. The index of refraction of the oil is 1.25 and that of the glass is 1.50. Find the wavelengths of light in the visible region (400 nm-750 nm) which are completely transmitted by the oil film under normal incidence.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Given that, λ = 560 * 10^–9 m, μ = 1.4. For strong reflection, 2μd = (2n + 1)λ/2 ⇒ d = (2n + 1)λ/4d For minimum thickness, putting n = 0. ⇒ d = λ/4d ⇒ d = 560 * 10^-9/14 = 10^-7 m = 100 nm.